# Improved Belief Propagation Decoding Algorithms for Surface Codes

### v3

@@ Abstract
Quantum error correction is crucial for universal fault-tolerant quantum computing. Highly accurate and low-time-complexity decoding algorithms play an indispensable role in ensuring quantum error correction works effectively. Among existing decoding algorithms, belief propagation (BP) is notable for its nearly linear time complexity and general applicability to stabilizer codes. However, BP’s decoding accuracy without post-processing is unsatisfactory in most situations. This article focuses on improving the decoding accuracy of BP over GF(4) for surface codes. We first propose Momentum-BP and AdaGrad-BP, inspired by machine learning optimization techniques, to reduce oscillation in message updating and break the symmetric trapping sets. We further propose EWAInit-BP, which adaptively updates initial probabilities and provides a 1 to 3 orders of magnitude improvement over traditional BP for planar surface code, toric code, and XZZX surface code without any post-processing method, showing high decoding accuracy even under parallel scheduling. The theoretical $O(1)$ time complexity under parallel scheduling and high accuracy of EWAInit-BP make it a promising candidate for high-precision real-time decoders. Meanwhile, the ideas of the Momentum-BP, AdaGrad-BP and EWAInit-BP provide promising approaches to improve the decoding accuracy of BP to get rid of its reliance on post-processing.
preprint: APS/123-QED
@@ I Introduction
Information in quantum computing devices is highly susceptible to noise, leading to information loss and computational errors [1]. To address this issue, quantum error correction (QEC) introduces redundant physical qubits and encodes the information in a certain subspace of the state space of all the physical qubits to protect it. So far at least, QEC is an essential step towards universal fault-tolerant quantum computation [2, 3, 4]. To ensure QEC works, the decoder of a QEC code should output an accurate and fast estimation of error according to the error syndromes.
Several decoding algorithms have achieved notable success, including minimum-weight perfect matching (MWPM) [5] algorithm, union-find (UF) algorithm [6], maximum likelihood (ML) algorithm [7], matrix-product-states-based (MPS-based) algorithm [7], renormalization group algorithm [8], and neural-network-based (NN-based) algorithm [9, 10, 11, 12, 13, 14, 15, 16]. Among these, ML, MWPM, and MPS-based algorithms offer high accuracy but come with high complexity. In contrast, the UF algorithm provides near-linear complexity with slightly lower accuracy. NN-based decoders have to balance scalability, decoding accuracy, and time cost [13], and may struggle with generalizability across different codes. Consequently, existing decoding algorithms rarely meet the requirements for both high accuracy and low complexity in quantum error correction decoding [17, 18].
Belief Propagation (BP) algorithm [19] are applicable to any stabilizer code and can achieve a complexity of $O(j\tau N)$ [20], where $j$ is the average weight of the rows and columns of the check matrix, and $\tau$ is the number of iterations. Additionally, parallel scheduling of BP can be efficiently implemented in hardware, theoretically achieving constant-time complexity [21]. However, traditional BP exhibits low decoding accuracy for many quantum error correcting codes and even cannot achieve the code capacity threshold for surface codes.
@@
In this article, we mainly focus on improving BP’s decoding accuracy for surface codes. However, we believe that the idea used in these improvements will also work for some other stabilizer codes. There are several explanations for why surface codes are challenging for BP decoding. From the perspective of degeneracy [37, 20], stabilizers of surface codes have much lower weights compared to the code distance, leading to multiple low-weight errors corresponding to the same syndrome, thus having negative influence on BP’s error convergence. From the perspective of short cycles [19], a number of 4-cycles over GF(4) and 8-cycles over GF(2) in surface codes can cause BP’s inaccurate updating. A more intuitive explanation involves trapping sets [38, 32], where surface codes have many symmetric trapping sets causing BP messages to oscillate periodically.
This article employs three strategies to improve BP: optimizing message updates by optimization methods in machine learning, and optimizing a priori probabilities by utilizing a posteriori probabilities output by previous iterations. First, inspired by the similarity between BP message update and gradient descent of the energy function, we leverage optimization methods in machine learning to optimize the message update rules, and propose Momentum-BP and AdaGrad-BP. In addition to improving decoding accuracy, experiments on trapping sets in surface codes show that Momentum-BP and AdaGrad-BP can smooth the message update process and reduce oscillation amplitude.
Furthermore, realizing the importance of a priori and a posteriori probabilities involved in the iteration of BP, we propose EWAInit-BP which dynamically updates the a priori probabilities involved in each iteration by utilizing the a posteriori probabilities output by previous iterations. Experiments on short toric codes, planar surface codes, and XZZX surface codes (under biased noise) show that EWAInit-BP outperforms existing BP improvements without post-processing and outer loops, under both parallel and serial scheduling. Notably, the time complexity of all proposed improvements remains $O(j\tau N)$. Thus, EWAInit-BP can be implemented with theoretically constant complexity in parallel while maintaining decoding accuracy.
@@ II.1 Surface codes
Surface codes may be the most notable stabilizer codes in recent years. Surface codes are a class of quantum Low-Density Parity-Check (LDPC) codes [39] where qubits are placed on a two-dimensional lattice, relying on the topology of the surface to protect quantum information [40]. Logical qubits are encoded in global degrees of freedom, and larger lattices provide larger code distances. There are some common properties of surface codes, such as the code distance scales with the code length as $O(\sqrt{N})$, while the code rate is asymptotically zero. In this article, toric codes refer to those with closed boundaries, and planar surface codes refer to those with open boundaries. The stabilizers of both types are either Pauli $X$ or Pauli $Z$. Besides, XZZX surface codes refer to the toric codes whose stabilizers are in the form of $XZZX$.
@@
A $[[2L^{2}-2L+1,1,L]]$ planar surface code is defined on a surface with open boundaries, which makes the stabilizers near the boundaries having lower weights. This reduction in stabilizer weight results in a slight decrease in the error correction capability compared to the toric code; however, it simplifies implementation in physical quantum systems.. The lattice structure of the planar surface codes are illustrated in Fig. 1.
@@
Figure 1: The lattice of the $[[2L^{2}-2L+1,1,L]]$ planar surface code with L=3. Circles represent data qubits, yellow squares represent X-type stabilizers, green squares represent Z-type stabilizers, and arrows indicate the coupling between stabilizers and qubits.
II.2 Belief propagation decoding
II.2.1 BP over GF(4)
@@
Besides, one can use log-likelihood ratio (LLR) as the message passed between variable nodes and check nodes. In this way, belief propagation (abbreviated as LLR-BP4) receives a priori probability vectors describing each type of Pauli error on each qubit $Q^{(0)}=\{p_{v_{j}}^{X},p_{v_{j}}^{Y},p_{v_{j}}^{Z}\}_{j=1}^{N}$, and initializes the variable-to-check messages as [20]
$m_{v_{j}\to c_{i}}^{(0)}=\lambda_{H_{ij}}(Q_{v_{j},e}^{(0)})=\lambda_{H_{ij}}(\ln\frac{1-p_{v_{j}}}{p_{v_{j}}}),$ (3)
where the lambda function is defined as
$\lambda(p^{X},p^{Y},p^{Z})=\ln\frac{1+e^{-p^{H_{ij}}}}{e^{-p^{X}}+e^{-p^{Y}}+e^{-p^{Z}}-e^{-p^{H_{ij}}}}$ (4)
indicating the log-likelihood ratio of whether the Pauli error $E_{j}$ commutes with the stabilizer $S_{i}$.
@@
$m_{c_{i}\to v_{j}}^{(t)}=(-1)^{s_{i}}*w.$ (6)
@@
In the vertical update, variable nodes aggregate the messages from check nodes and, after a similar extrinsic calculation as above, send variable-to-check messages as follows
@@
$m_{v_{j}\to c_{i}}^{(t)}=\lambda_{H_{ij}}(Q_{v_{j},e}^{(0)}+\sum_{\begin{subarray}{c}c_{i^{\prime}}\in\mathcal{M}(v_{j})\backslash c_{i}\\ \left\langle e,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i^{\prime}}\rightarrow v_{j}}^{(t)}).$ (7)
@@
In the hard decision, variable nodes first accumulate the messages from all check nodes to calculate the a posteriori probability
@@
$Q_{v_{j},e}^{(t)}=Q_{v_{j},e}^{(0)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle e,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)}.$ (8)
@@
Then, for each variable node, if all log-likelihood ratios for Pauli errors are greater than 0, then the hard decision result is the identity $I$; otherwise, it is the Pauli error with the smallest LLR value. The hard decisions of all qubits combine to form the error estimate $\hat{E}$. If $\hat{E}$ corrects the error syndrome, the algorithm converges; otherwise, it returns to the horizontal update for a new iteration. If the maximum number of iterations is reached, decoding is considered to have failed.
@@
BP can be performed under either parallel or serial scheduling [42]. Under parallel scheduling, all variable nodes and check nodes update their messages simultaneously in each iteration, allowing for hardware parallelization and constant time complexity. On the other hand, the serial scheduling updates the messages of variable nodes and check nodes sequentially. This method leads to faster convergence and slightly higher accuracy, though it has a more complex hardware implementation and higher time complexity.
II.2.2 BP as gradient optimization
$J_{S}(\Gamma)=$ $-\sum_{i=1}^{m}2\tanh^{-1}$
$\left((-1)^{s_{i}}\prod_{v_{j}\in\mathcal{N}(c_{i})}\tanh\left(\frac{\lambda_{H_{ij}}(Q_{v_{j}})}{2}\right)\right).$ (9)
The term inside the parentheses of the energy function $J_{S}$ measures the discrepancy between the result of the current iteration and the actual error syndrome of stabilizer $S_{i}$. The smaller the energy function, the closer the BP iteration result is to the correct solution.
The target variables for gradient updates are the a posteriori estimates of the probabilities of each type of Pauli error occurring at each variable node. Specifically, for the $e\in\{X,Y,Z\}$ type Pauli error on the $n$-th variable node, the a posteriori probability update is
$Q_{v_{j},e}^{(t)}=Q_{v_{j},e}^{(t-1)}-\alpha\frac{\partial J}{\partial Q_{v_{j},e}^{(t-1)}},$ (10)
where the partial derivative is
$\frac{\partial J}{\partial Q_{v_{j}}^{e}}=$ $-\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle e,H_{ij}\right\rangle=1\end{subarray}}\widetilde{m}_{c_{i}\rightarrow v_{j}}$
$*\frac{\eta g_{ij}(\Gamma)e^{-Q_{v_{j}}^{e}}}{e^{-Q_{v_{j}}^{X}}+e^{-Q_{v_{j}}^{Y}}+e^{-Q_{v_{j}}^{Z}}-e^{-Q_{v_{j}}^{H_{ij}}}},$ (11)
where $\widetilde{m}_{c_{i}\rightarrow v_{j}}$ is the a posteriori probability summation term, approximated using the $\tanh$ function
$\widetilde{m}_{c_{i}\rightarrow v_{j}}=(-1)^{s_{i}}\prod_{v_{j^{\prime}}\in\mathcal{N}(c_{i})\backslash v_{j}}\tanh\left(\frac{\lambda_{H_{ij^{\prime}}}(\Gamma_{v_{j^{\prime}}})}{2}\right).$ (12)
@@ II.2.2 BP as gradient optimization
From this gradient optimization perspective, MBP [20] adjusts the ”learning rate” of message updates to converge more quickly to the equivalent energy minimum, which is common in degenerate quantum codes.
II.2.3 Hardness of BP in decoding surface codes
The decoding accuracy of BP in decoding surface codes is unsatisfactory due to several reasons, including cycles, degeneracy, and trapping sets.
@@ II.2.3 Hardness of BP in decoding surface codes
Short cycles cause the joint probability calculations in BP to be dependent, which violates the independence assumption of updating variable nodes Eq. (7) and check nodes Eq. (5) [19]. Surface codes exhibit 8-cycles over GF(2) and 4-cycles over GF(4), exacerbating this issue.
@@
High degeneracy leads to BP decoding objectives Eq. (2) deviating from true targets Eq. (1) [22, 37]. This problem becomes particularly severe when errors can combine with stabilizers to form error types that are symmetric to the original ones. In such cases, BP’s marginal probability calculations for the both error types will be completely consistent. In surface codes, the weights of the stabilizers are all no greater than 4, often much less than the weights of the logical operators. This symmetric degeneracy results in incorrect convergence of BP decoding.
@@
Trapping sets are specific local structures in error-correcting codes combined with corresponding error patterns, where iterative decoding algorithms can never converge [44, 45, 46]. In quantum error correction, trapping sets are caused by the aforementioned short loops and symmetric degeneracy [38, 47]. The smallest trapping set in surface codes is shown in Fig. 2, consisting of an 8-cycle and four symmetric stabilizers, where the local connections of each check node and variable node are identical [32]. If any two qubits in this trapping set experience errors that anticommute with the stabilizers, the hard decision result of BP will oscillate periodically between all 0s and all 1s.
Optimizing BP from the first two perspectives involves modifying the underlying principles of the algorithm. This article will focus on trapping sets and seek methods to break the symmetry in message passing.

### v5

@@ Abstract
Quantum error correction is crucial for universal fault-tolerant quantum computing. Highly accurate and low-time-complexity decoding algorithms play an indispensable role in ensuring quantum error correction works effectively. Among existing decoding algorithms, belief propagation (BP) is notable for its nearly linear time complexity and general applicability to stabilizer codes. However, BP’s decoding accuracy without post-processing is unsatisfactory in most situations. This article focuses on improving the decoding accuracy of BP over GF(4) for surface codes. Inspired by machine learning optimization techniques, we first propose Momentum-BP and AdaGrad-BP to reduce oscillations in message updating, breaking the trapping sets of surface codes. We further propose EWAInit-BP, which adaptively updates initial probabilities and provides a 1 to 3 orders of magnitude improvement over traditional BP for planar surface code, toric code, and XZZX surface code without any post-processing method, showing high decoding accuracy even under parallel scheduling. The theoretical $O(1)$ time complexity under parallel implementation and high accuracy of EWAInit-BP make it a promising candidate for high-precision real-time decoders.
Index Terms:
Quantum Error Correction, Quantum Information, Machine Learning
[…]
=-15pt
@@ I Introduction
Information in quantum computing devices is highly susceptible to noise, leading to information loss and computational errors [1]. To address this issue, quantum error correction (QEC) introduces redundant physical qubits and encodes the information in a certain subspace of the state space of all the physical qubits to protect it. So far at least, QEC is an essential step towards universal fault-tolerant quantum computation [2, 3, 4]. To ensure QEC works, the decoder of a QEC code should output an accurate and fast estimation of error according to the error syndromes [5].
Several decoding algorithms have achieved notable success, including minimum-weight perfect matching (MWPM) [6] algorithm, union-find (UF) algorithm [7], maximum likelihood (ML) algorithm [8], matrix-product-states-based (MPS-based) algorithm [8], renormalization group algorithm [9], and neural-network-based (NN-based) algorithm [10, 11, 12, 13, 14, 15, 16, 17]. Among these, ML, MWPM, and MPS-based algorithms offer high accuracy but come with high complexity. In contrast, the UF algorithm provides near-linear complexity with slightly lower accuracy. NN-based decoders have to balance scalability, decoding accuracy, and time cost [14], and may struggle with generalizability across different codes. Consequently, existing decoding algorithms rarely meet the requirements of both high accuracy and low complexity in quantum error correction decoding [18, 19].
Belief Propagation (BP) algorithm [20] are applicable to any stabilizer code and can achieve a complexity of $O(j\tau N)$ [21], where $j$ is the average weight of the rows and columns of the check matrix, and $\tau$ is the number of iterations. Additionally, parallel scheduling of BP can be efficiently implemented in hardware, theoretically achieving constant-time complexity [22]. However, traditional BP exhibits low decoding accuracy for many quantum error correcting codes and even cannot achieve the code capacity threshold for surface codes, which are the focus of recent advancements in quantum error correction implementation [3, 23].
@@
In this article, we mainly focus on improving BP’s decoding accuracy for surface codes, while we also believe that the ideas used in these improvements will also work for some other quantum LDPC codes. There are several explanations for why surface codes are challenging for BP decoding. From the perspective of degeneracy [39, 21], stabilizers of surface codes have much lower weights compared to the code distance, leading to multiple low-weight errors corresponding to the same syndrome, thus having negative influence on BP’s error convergence. From the perspective of short cycles [20], a number of 4-cycles over GF(4) and 8-cycles over GF(2) in surface codes can cause BP’s inaccurate updating. A more intuitive explanation involves trapping sets [40, 34], where surface codes have many symmetric trapping sets causing BP messages to oscillate periodically.
This paper approaches BP improvements from two perspectives: optimizing message updates inspired by gradient optimization, and optimizing a priori probabilities using information of previous iterations, leading to the proposal of three algorithms.
First, inspired by the similarity between BP message update and gradient descent of the energy function, we leverage optimization methods in machine learning to optimize the message update rules, and propose Momentum-BP and AdaGrad-BP. In addition to improving decoding accuracy, experiments on trapping sets in surface codes show that Momentum-BP and AdaGrad-BP can smooth the message update process and reduce oscillation amplitude.
Furthermore, realizing the importance of a priori probabilities involved in the iteration of BP, we propose EWAInit-BP which dynamically updates the a priori probabilities involved in each iteration by utilizing the a posteriori probabilities output by previous iterations. Experiments on short toric codes, planar surface codes, and XZZX surface codes (under biased noise) show that EWAInit-BP outperforms existing BP improvements without post-processing and outer loops under parallel scheduling. Notably, the time complexity of all proposed improvements remains $O(j\tau N)$. Thus, EWAInit-BP can be implemented with theoretically constant complexity in parallel while maintaining decoding accuracy.
@@ II-A Surface codes
Surface codes may be the most notable stabilizer codes in recent years. Surface codes are a class of quantum Low-Density Parity-Check (LDPC) codes [41] where qubits are placed on a two-dimensional lattice, relying on the topology of the surface to protect quantum information [42]. Logical qubits are encoded in global degrees of freedom, and larger lattices provide larger code distances. The code distance scales with the code length as $O(\sqrt{N})$, whereas the code rate is asymptotically zero. In this article, toric codes refer to those with closed boundaries, and planar surface codes refer to those with open boundaries. Besides, We consider the unrotated surface codes, but it is clear that our proposed algorithms can be extended to rotated versions.
@@
A $[[2L^{2}-2L+1,1,L]]$ planar surface code is defined on a surface with open boundaries, which makes the stabilizers near the boundaries having lower weights. This reduction in stabilizer weight results in a slight decrease in the error correction capability compared to the toric code; however, it simplifies implementation in physical quantum systems.. The lattice structure of the planar surface codes are illustrated in Fig. II-A.
@@
[…]
[t!](topskip=-10pt, botskip=0pt, midskip=0pt)[width=2.5in]chen1.pdf The lattice of the $[[2L^{2}-2L+1,1,L]]$ planar surface code with L=3. Circles represent data qubits, yellow squares represent X-type stabilizers, green squares represent Z-type stabilizers, and arrows indicate the coupling between stabilizers and qubits.
II-B Belief propagation decoding
II-B1 BP over GF(4)
@@
Besides, one can use log-likelihood ratio (LLR) as the message passed between variable nodes and check nodes. In this way, belief propagation (abbreviated as LLR-BP4) receives a priori probability vectors describing each type of Pauli error on each qubit as $\{p_{v_{j}}=\{p_{v_{j}}^{X},p_{v_{j}}^{Y},p_{v_{j}}^{Z}\}\}_{j=1}^{N}$, and initializes the variable-to-check messages as [21]
$\Pi_{v_{j}}$ =$=\{{\Pi_{v_{j},W}}\}_{W\in\{X,Y,Z\}}=\ln\frac{1-p_{v_{j}}}{p_{v_{j}}},$ (3)
$m_{v_{j}\to c_{i}}^{(0)}$ $=\lambda_{H_{ij}}({\Pi_{v_{j},X}},{\Pi_{v_{j},Y}},{\Pi_{v_{j},Z}}),$ (4)
where $\Pi_{v_{j}}$ is the a priori probabilities of variable node $j$ expressed by LLR. The lambda function in (4) is defined as
$\lambda_{H_{ij}}(p^{X},p^{Y},p^{Z})=\ln\frac{1+e^{-p^{H_{ij}}}}{e^{-p^{X}}+e^{-p^{Y}}+e^{-p^{Z}}-e^{-p^{H_{ij}}}}$ (5)
indicating LLR values of whether the Pauli error $E_{j}$ commutes with the Pauli operator $H_{ij}$ of the parity-check matrix $H$. $H_{ij}$ is also the $j$-th element of stabilizer $S_{i}$.
@@
$m_{c_{i}\to v_{j}}^{(t)}=(-1)^{s_{i}}*w,$ (7)
where $\mathcal{N}(c_{i})$ represents all variable nodes adjacent to check node $i$, and $s_{i}$ is the syndrome corresponding to stabilizer $S_{i}$.
@@
In the vertical update, variable nodes aggregate the messages from check nodes and, after a similar extrinsic calculation as above, send variable-to-check messages as
@@
$m_{v_{j}\to c_{i}}^{(t)}=\lambda_{H_{ij}}(\{\Pi_{v_{j},W}+\sum_{\begin{subarray}{c}c_{i^{\prime}}\in\mathcal{M}(v_{j})\backslash c_{i}\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i^{\prime}}\rightarrow v_{j}}^{(t)}\}_{W\in\{X,Y,Z\}}),$ (8)
where $\mathcal{M}(v_{j})$ represents all check nodes adjacent to variable node $j$. $\left\langle W,H_{ij}\right\rangle$ stands for the trace inner product, here the result is 1 when the error operator $W$ anti-commutes with the matrix element $H_{ij}$ and 0 when they commute.
@@
In the hard decision, variable nodes first accumulate the messages from all check nodes to calculate the a posteriori probability of each Pauli error on each qubit as
@@
$Q_{v_{j},W}^{(t)}=\Pi_{v_{j},W}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)}.$ (9)
@@
Then, for each variable node, if all log-likelihood ratios for Pauli errors are greater than 0, then the hard decision result is the identity $I$; otherwise, it is the Pauli error with the smallest LLR value. The hard decisions of all qubits combine to form the error estimate $\hat{E}$. If $\hat{E}$ corrects the error syndrome, the algorithm converges; otherwise, it returns to the horizontal update for a new iteration. If the maximum number of iterations is reached, decoding has failed.
@@
BP can be performed under either parallel or serial scheduling [44]. Under parallel scheduling, all variable nodes and check nodes update their messages simultaneously in each iteration, allowing for hardware parallelization and constant time complexity. On the other hand, the serial scheduling updates the messages of variable nodes and check nodes sequentially. This method leads to faster convergence and slightly higher accuracy, though it has a more complex hardware implementation and higher computational overhead.
@@
One perspective on BP is that each a posteriori probabilities update in one BP iteration can be viewed as a single step of gradient descent with a simplified step size on a certain energy function, as the expanded form of the a posteriori update closely resembles the derivative of the energy function (see Appendix VI-A). From this perspective, MBP and AMBP [21] adjusts the ”learning rate” of message updates to converge more quickly to the equivalent energy minimum, which is common in degenerate quantum codes.
II-B2 Hardness of BP in decoding surface codes
There are several terms that can explain the unsatisfactory accuracy of BP in surface codes.
@@ II-B2 Hardness of BP in decoding surface codes
Short cycles cause the joint probability calculations in BP to be dependent, which violates the independence assumption of updating variable nodes (8) and check nodes (6) [20]. Surface codes exhibit 8-cycles over GF(2) and 4-cycles over GF(4), exacerbating this issue.
@@
High degeneracy leads to BP decoding objectives (2) deviating from true targets (1) [24, 39], especially when errors can combine with stabilizers to form error types that are symmetric to the original ones. In such cases, BP’s marginal probability calculations for the both error types will be completely consistent. In surface codes, the weights of the stabilizers are all no greater than 4, often much less than the weights of the logical operators. This symmetric degeneracy results in incorrect convergence of BP decoding.
@@
Trapping sets are specific local structures in error-correcting codes combined with corresponding error patterns, where iterative decoding algorithms can never converge [45, 46, 47]. In QEC, trapping sets are caused by the aforementioned short loops and symmetric degeneracy [40, 48]. The smallest trapping set in surface codes is shown in Fig. 1, consisting of an 8-cycle and four symmetric stabilizers, where the local connections of each check node and variable node are identical [34]. If any two qubits here experience errors that anticommute with the stabilizers, the hard decision result of BP will oscillate periodically between all 0s and all 1s.
The first two terms are directly related to the underlying principles of the BP algorithm. In contrast, trapping sets offer a more specific description, which can not only inspire potential improvements but also serve as a metric for evaluating improvements.